Solve each of the following systems by substitution.
8x - 4y = 16
y = 2x - 4

Answer by praseenakos@yahoo.com(507)

Question:
Solve each of the following systems by substitution.
8x - 4y = 16
y = 2x - 4

Answer:
Given equations are,
8x - 4y = 16
y = 2x - 4

substitute the value y = 2x-4 in the first equation...

8x - 4(2x - 4) = 16

==> 8x - 4*2x - 4*4 = 16

==> 8x - 8x - 16 = 16
==> -16 = 16, which is not a true statement...

So the given system equation has no solution..


Hope you found the explanation useful.

Regards.

Praseena.